January 1,2021, Raicel Co. traded an old equipment for a newer model. Data before the trade in are follows:

Old equipment:

Cost ₱350,000

Accumulated depreciation 120,000

Average published retail value 60,000

New equipment:

List price ₱400,000

Cash price without trade in 340,000

Cash price with trade in 270,000

Questions:

Based on the above data, answer the following:

1. How much is the cost of the equipment to be recognized by Raciel Co.?

2. How much is the gain (or loss) on trade in to be recognized by Raciel Co.?



1
Expert's answer
2021-06-21T11:58:25-0400

Solution:

1.). The cost of the equipment to be recognized by Raciel Co. is ₱180,000

Loss or gain = (Fair value – cash price with trade in) – (Original value of old equipment – accumulated depreciation)

= (340,000 – 270,000) – (350,000 – 120,000) = 70,000 – 230,000 = (160,000)

Loss on trade in = (160,000)

Cost of the equipment = (Original value of old equipment – accumulated depreciation) + cash price with trade-in – loss on trade-in

Cost of the equipment = (350,000 – 120,000) + 270,000 – 160,000 = 180,000

Cost of the equipment to be recognized by Raciel Co. = ₱180,000

 

2.). The gain or loss on trade in to be recognized by Raciel Co. is a loss of (₱160,000)

Loss or gain = (Fair value – cash price with trade in) – (Original value of old equipment – accumulated depreciation)

= (340,000 – 270,000) – (350,000 – 120,000) = 70,000 – 230,000 = (160,000)

Loss on trade in = (₱160,000)